Suppose that in the short run, the wage rate is $10 per worker and capital costs are fixed at $100. If the firm employs 5 workers, it will be able to produce 20 units of output. What is the average total cost of producing 20 units of output?

  1. $7.50
  2. $4
  3. $2.50
  4. $0.40
  5. None of the above
Answer #1

Ans) the correct option is a) $ 7.50

Total cost = 5*10 + 100 = 150

Average total cost = 150/20 = 7.50